OpenStudy (hamzajani):

The force vector is F=2i+6j+7k and body covers a displacement of 2m along y-axis the work done is.....? A)√89×2j B)3i+6kj C)10j D)13j

OpenStudy (sharmagk83):

Data Given. force \[F = 2i + 6 j + 7k\] and displacement \[s = 2 j\] Work done is given by dot product of force vector and displacement vector. So \[W = F.s = (2i+6j+7k).(2j) = 4 i.j +12 j.j +14 k.j = 4\times0 +12\times1+14\times0\] \[W = 12 J\]